It has been found that on average, four motorbikes go through an N1 tollgate every 20 minutes. What is the probability that

No more than 2 motorbikes will go through the tollgate in any randomly selected 20-minute period? Interpret your answer.


1
Expert's answer
2021-10-25T03:13:11-0400

Let "X=" the number of motorbikes through the tollgate: "X\sim Po(\lambda t)"

Given "\lambda t=4."


"P(X\leq 2)=P(X=0)+P(X=1)+P(X=2)"

"=\dfrac{e^{-4}\cdot4^0}{0!}+\dfrac{e^{-4}\cdot4^1}{1!}+\dfrac{e^{-4}\cdot4^2}{2!}=13e^{-4}"




"\approx0.238103"

The probability that no more than 2 motorbikes will go through the tollgate in any randomly selected 20-minute period is "0.238103."